OpenStudy (anonymous):

Write the equation of the line that passes through (2, 2) and (6, 3) in standard form.

OpenStudy (anonymous):

y = mx + b 2 = 2m + b 3 = 6m + b 6 = 6m + 3b - 3 = 6m + b ------------- 3 = 2b 1.5 = b 2 = 2m + b 2 = 2m + 1.5 0.5 = 2m m = 0.25 y = 0.25x + 1.5

OpenStudy (anonymous):

It's none of them. Are you sure you wrote down the question right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

4x + 4y = 6 x + 4y = 12 4x + 4y = –12 x – 4y = –6

OpenStudy (anonymous):

Well, then there's something wrong with the question. Just try substituting for x and y. When x = 2 , y=2, and when x = 6 , y = 6.

OpenStudy (anonymous):

i did a mistake thats the answers 4x + 4y = 6 x + 4y = 12 4x + 4y = –12 x – 4y = –6